# CogFlow: Bridging Perception and Reasoning through Knowledge Internalization for Visual Mathematical Problem Solving

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) are rapidly advancing and have been applied across various vision–language applications (Peng et al., 2024b; Xu et al., 2025b; Yue et al., 2024). However, existing MLLMs continue to struggle with challenging visual mathematical problems, resulting in low answer accuracy and inconsistent reasoning chains. Some early attempts (Wang et al., 2025a; Shen et al., 2025) adopt a one-step reasoning framework that directly interleaves visual perception with reasoning in an unstructured manner, often resulting in both perceptual and reasoning errors. Another line of work (Chen et al., 2025a; Guo et al., 2025c) follows a decoupled reasoning pipeline that explicitly separates the perception and reasoning parts, with the former focusing on visual recognition and the latter responsible for subsequent inference. Yet in practice, we observe that such a pipeline often suffers from the reasoning drift issue, i.e., it tends to yield illogical or unwarranted reasoning steps that disregard perceptual evidence (see Figure 1). These observations motivate the development of a new approach that not only achieves robust fine-grained recognition of mathematical visual elements (e.g., diagrammatic primitives and symbols) but also faithfully incorporates extracted visual cues into subsequent reasoning.
@@
Unlike prior approaches (Jia et al., 2024; Chen et al., 2025a; Guo et al., 2025c) that decouple perception from reasoning trajectories (Ouyang et al., 2022) and enhance it with tailored tasks, CogFlow first integrates perception enhancement into a unified reinforcement learning (RL) framework through Synergistic Visual Rewards (SynVRs), enabling dynamic perception–reasoning interaction and improving generalization. Specifically, SynVRs complementarily optimize the model from two distinct perspectives: (1) a Visual Parameterized Reward (VPR) encoding normalized primitives (i.e., points, lines, and circles) and calculating the Euclidean distance in a parameter space for precise and interpretable measurement; (2) a Visual Semantic Reward (VSR) that extracts semantic embeddings (Xie et al., 2025) from re-rendered images (derived from textual perception outputs) and measures the cosine distance in a semantic space to capture holistic style and layout consistency. Together, SynVRs ensure both local geometric fidelity and global perceptual coherence, forming trustworthy visual cues that serve as the foundation for effective visual mathematical reasoning.
Notably, despite progress in perception enhancement, all prior efforts (Jia et al., 2024; Guo et al., 2025c; Wei et al., 2025) remain confined to accurate extraction of mathematical information from diagrams, while ignoring a key question: are the extracted visual cues properly and faithfully integrated into subsequent reasoning? As illustrated in Figure 1(b), our empirical findings reveal a typical reasoning drift issue (i.e., the reasoning stage in existing methods often deviates from perceptual results), leading to reasoning chains that appear coherent yet conflict with the underlying visual evidence. To prevent such drift and improve interpretability, CogFlow utilizes a Knowledge Internalization Reward (IntlzR) that bridges the perception and reasoning stages by encouraging the model to generate structured and reasoning-ready outputs (i.e., knowledge-internalized representations (Ryan and Connell, 1989)) as a more reliable foundation for subsequent reasoning. Specifically, we curate positive trajectories integrating perception and reasoning processes with explicit internalization of perception primitives, and further derive five typical negative trajectories. Training with these trajectories enables the reward model to evaluate each response according to its fidelity to the internalized representation. IntlzR effectively improves the knowledge internalization stage, thereby reducing hallucinations and improving interpretability and robustness.
@@ 2 Related Work
Visual Mathematical Reasoning with MLLMs. Solving visual mathematical problems (e.g., geometry diagrams, algebraic plots, and etc) requires both strong reasoning ability and accurate interpretation of visual primitives and symbolic content (Yan et al., 2025; Qiao et al., 2025b; Zhang et al., 2025b). Most previous works are dedicated to improving the reasoning process, including chain-of-thought strategies (Xu et al., 2025a; Deng et al., 2024), tool-aided reasoning (Trinh et al., 2024; Chen et al., 2023b), test time scaling (Wang et al., 2025e; Hosseini et al., 2024), and reinforcement learning (Wang et al., 2025c; Jiang et al., 2025). Several recent works (Guo et al., 2025c; Jia et al., 2024; Wei et al., 2024) suggest that one of the major bottlenecks in visual mathematical reasoning is inaccurate visual comprehension. They typically decouple perception from reasoning, and strengthen perception either by designing specialized visual encoders (Zhang et al., 2025a) or by introducing auxiliary visual tasks (Chen et al., 2025a). However, prior works ignore a key issue of whether correctly extracted visual cues are indeed faithfully incorporated into subsequent reasoning.
@@ 3.1 Forming Trustworthy Perception with Synergistic Visual Rewards
We then compute the cost matrix $\mathcal{C}$ between GT Primitives $\mathcal{G}$ and predicted Primitives $\mathcal{P}$ and apply the Hungarian matching algorithm (Kuhn, 1955) to obtain the optimal one-to-one matching $\mathcal{H}$ that minimizes the total cost $\mathcal{S}_{\text{VPR}}$. The VPR offers interpretable, geometry-aware semantic supervision that avoids the pitfalls of pixel-level noise or black-box embedding similarity. More details are provided in §C.2.
@@ 3.2 Alleviating Reasoning Drift with Knowledge Internalization Reward
Rewarding Faithful Knowledge Internalization. To address these systematic weaknesses, we introduce the Knowledge Internalization Reward (IntlzR), a trained reward model that evaluates whether each reasoning chain remains faithful to the internalized visual representation, thereby enforcing raw perception as the primary substrate of inference (e.g., reasoning drift). Within CogFlow, we train the IntlzR to enhance internalization explicitly, supported by the curated MathCog-IntlzR (see §C.3 for more details). This dataset is constructed from positive–negative pairs, with each positive matched to five corresponding negatives; positives are sampled from MathCog-SFT, and negatives are synthesized by injecting the five error types described above, providing fine-grained supervision for distinguishing grounded from unanchored reasoning. To better leverage diverse negative signals in IntlzR and adaptively emphasize the most challenging trajectories while improving gradient efficiency and training stability, we adopt Softmax-DPO (Chen et al., 2024a) for optimization:
@@ 3.3 Strengthening Multi-Step Visual Reasoning with VGPO
Then the training optimization objective $\mathcal{L}$ is formulated as:
@@ 4.1 Experimental Setup
Baselines. We compare our method with a wide range of closed-source MLLMs, such as GPT series (OpenAI, 2024; 2023; 2025), Gemini series (Team et al., 2025), DouBao-1.5-pro (Guo et al., 2025b), GLM-4.5V (Hong et al., 2025), Claude (Anthropic, 2024) and Qwen-VL-Plus (Bai et al., 2023), as well as open-source MLLMs, including InfiMM-Math (Han et al., 2025), InternVL2.5 (Chen et al., 2024b), InternLM-XC2 (Dong et al., 2024), Math-LLaVA (Shi et al., 2024), MultiMath-7B (Peng et al., 2024a), SVE-Math (Zhang et al., 2025b), VLM-R1 (Shen et al., 2025), MathFlow (Chen et al., 2025a), DVLR (Guo et al., 2025c), VCAR (Jia et al., 2024), VL-Rethinker (Wang et al., 2025a), ThinkLite-VL (Wang et al., 2025d) and SPHINX-MoE (Liu et al., 2024a). Since some concurrent works have yet to release their models, we report the results provided in their original papers, some of which do not include category-wise performance.
Implementation Details. We initialize CogFlow with Qwen2.5-VL-7B (Bai et al., 2025) and train it on our curated MathCog dataset. For cold start, we first train CogFlow on MathCog-SFT subset for 2 epochs with a learning rate of $1\times 10^{-5}$ and a batch size of 64. Subsequently, CogFlow is optimized using VGPO on MathCog-RL subset for 1 epoch with a learning rate of $1\times 10^{-6}$ and a batch size of 16. The IntlzR reward model is based on Qwen2.5-VL-3B and trained on MathCog-IntlzR subset for 3 epochs with a learning rate of $7\times 10^{-6}$ and a batch size 64. All our models are trained on 16 NVIDIA A100 GPUs. Please refer to §C for more detailed configures.
@@ 4.2 Main Results
As shown in Tables 14.2, CogFlow achieves 66.0% accuracy on FlowVerse, 53.9% on MathVerse, 76.8% on Mathvista, 64.1% accuracy on WeMath and 46.2% on DynaMath, surpassing all open-source baselines by large margins. In terms of LogicVista, it delivers competitive performance, trailing only GLM-4.IV-9B. CogFlow’s improvements in visual perception are particularly significant, as evidenced by its superior performance on subsets where visual information dominates. For example, on FlowVerse it reaches 42.7% on Vision Dense, 55.6% on Vision Primary, and 61.1% on Vision Centric. On MathVerse, the gains are consistent across Vision Intensive (44.8%), Vision Dominant (42.1%), and Vision Only (25.7%). These results demonstrate that CogFlow yields more accurate parsing of geometric primitives and relations, leading to stronger perception–reasoning integration. CogFlow also delivers higher-end-to-end problem-solving accuracy across all settings. The largest absolute gains appear in visually demanding subsets, where perception-anchored reasoning is essential. This demonstrates that CogFlow not only perceives diagrams more accurately but also reasons over them more effectively. Figure 19 further shows that CogFlow produces more structurally accurate and semantically consistent predictions compared to strong baselines. Please refer to §D.12 for more detailed configurations.
@@ 4.3 Ablation Studies
Analysis on Knowledge Internalization Reward. As shown in Figure 6, removing any single Error type consistently degrades performance, indicating that each contributes complementary supervision for internalization. The largest drops arise when excluding omission/misbinding primitives or contradicting geometric constraints, highlighting that correctly binding primitives and respecting core geometric constraints are most critical for keeping reasoning tied to perception. Figure 6 compares different training strategies during training. Vanilla DPO with all categories offers moderate gains, whereas Softmax-DPO achieves the best results, reaching 66.0% / 56.2% on FlowVerse. This suggests that fine-grained, within-problem preference modeling better captures the nuances of perception-grounded reasoning.
@@
Error Type Analysis. As shown in Figure 9, we employ GPT-5 (OpenAI, 2025) to classify each response on the FlowVerse benchmark into one of four categories: Perception Error, Knowledge Internalization Error, Reasoning Error, and Correct. The first row covers specialized visual-math MLLMs (MultiMath-7B (Peng et al., 2024a), SVE-Math-7B (Zhang et al., 2025b)), an GRPO-style model (VLM-R1 (Shen et al., 2025)), the decoupled method (MathFlow-7B (Chen et al., 2025a)), and GPT-4o (OpenAI, 2024); the second row traces our variants from the SFT+GRPO baseline through +SynVRs, +IntlzR, and +VGPO to CogFlow. We observe: (1) GPT-4o’s strong overall performance is driven by superior reasoning, yet its diagram perception lags behind specialized 7B models, indicating that strengthening perception and internalization remains essential; (2) Although methods such as MathFlow-7B and VLM-R1 reduce Perception Error and Reasoning Error to some extent, these methods do not meaningfully alleviate Knowledge Internalization Error; (3) +SynVRs chiefly reduces Perception Errors ($2\%$), confirming that geometry-aware and semantic visual rewards improve perceptual fidelity; (4) +IntlzR primarily mitigates reasoning drift, reducing Knowledge-Internalization Error by $2\%$ relative to the baseline; and (5) when all components are combined, CogFlow minimizes every error type while maximizing the proportion of correct predictions. Furthermore, we conduct a systematic case study (see §D.13) to analyze the error compensation mechanism.
@@ B.4 Dataset Examples
Specifically, the perception trajectory translates raw diagrammatic pixels into structured geometric primitives (Points, Lines, and Circles) with precise coordinates in a visual parameterized space. For instance, in the leftmost example involving a rectangle, the model identifies specific coordinates for vertices $A$ and $B$ (e.g., A:(7.56,3.5)A:(-7.56,3.5)) and the intersection point $O$. Similarly, in the center and right examples, the model explicitly parameterizes circles by identifying center coordinates and radii (e.g., Circle: (2.52,0.48,7.84)(2.52,-0.48,7.84)).
The reasoning trajectory section then performs logical inference anchored in these extracted visual cues. For example, in the rightmost case, the reasoning process explicitly links the perceived fact that ”AB is the diameter” (derived from the coordinates in <WATCHING>) to the geometric theorem that the angle subtended by a diameter is $90^{\circ}$, leading to the correct deduction that $\angle ACB=90^{\circ}$5. This structured format ensures that the final <ANSWER> is derived from a cognitive flow of perception $\Rightarrow$ internalization $\Rightarrow$ reasoning.

### v3

@@ 1 Introduction
Multimodal large language models (MLLMs) are rapidly advancing and have been applied across various vision–language applications (Peng et al., 2024b; Xu et al., 2024; 2025b; Yue et al., 2024; Liang et al., 2025; Jia et al., 2024a; Quan et al., 2024; Zhang et al., 2025c; Ma et al., 2024a; b). However, existing MLLMs continue to struggle with challenging visual mathematical problems, resulting in low answer accuracy and inconsistent reasoning chains. Some early attempts (Wang et al., 2025a; Shen et al., 2025) adopt a one-step reasoning framework that directly interleaves visual perception with reasoning in an unstructured manner, often resulting in both perceptual and reasoning errors. Another line of work (Chen et al., 2025a; Guo et al., 2025c) follows a decoupled reasoning pipeline that explicitly separates the perception and reasoning parts, with the former focusing on visual recognition and the latter responsible for subsequent inference. Yet in practice, we observe that such a pipeline often suffers from the reasoning drift issue, i.e., it tends to yield illogical or unwarranted reasoning steps that disregard perceptual evidence (see Figure 1). These observations motivate the development of a new approach that not only achieves robust fine-grained recognition of mathematical visual elements (e.g., diagrammatic primitives and symbols) but also faithfully incorporates extracted visual cues into subsequent reasoning.
@@
Unlike prior approaches (Jia et al., 2024b; Chen et al., 2025a; Guo et al., 2025c) that decouple perception from reasoning trajectories (Ouyang et al., 2022) and enhance it with tailored tasks, CogFlow first integrates perception enhancement into a unified reinforcement learning (RL) framework through Synergistic Visual Rewards (SynVRs), enabling dynamic perception–reasoning interaction and improving generalization. Specifically, SynVRs complementarily optimize the model from two distinct perspectives: (1) a Visual Parameterized Reward (VPR) encoding normalized primitives (i.e., points, lines, and circles) and calculating the Euclidean distance in a parameter space for precise and interpretable measurement; (2) a Visual Semantic Reward (VSR) that extracts semantic embeddings (Xie et al., 2025) from re-rendered images (derived from textual perception outputs) and measures the cosine distance in a semantic space to capture holistic style and layout consistency. Together, SynVRs ensure both local geometric fidelity and global perceptual coherence, forming trustworthy visual cues that serve as the foundation for effective visual mathematical reasoning.
Notably, despite progress in perception enhancement, all prior efforts (Jia et al., 2024b; Guo et al., 2025c; Wei et al., 2025) remain confined to accurate extraction of mathematical information from diagrams, while ignoring a key question: are the extracted visual cues properly and faithfully integrated into subsequent reasoning? As illustrated in Figure 1(b), our empirical findings reveal a typical reasoning drift issue (i.e., the reasoning stage in existing methods often deviates from perceptual results), leading to reasoning chains that appear coherent yet conflict with the underlying visual evidence. To prevent such drift and improve interpretability, CogFlow utilizes a Knowledge Internalization Reward (IntlzR) that bridges the perception and reasoning stages by encouraging the model to generate structured and reasoning-ready outputs (i.e., knowledge-internalized representations (Ryan and Connell, 1989)) as a more reliable foundation for subsequent reasoning. Specifically, we curate positive trajectories integrating perception and reasoning processes with explicit internalization of perception primitives, and further derive five typical negative trajectories. Training with these trajectories enables the reward model to evaluate each response according to its fidelity to the internalized representation. IntlzR effectively improves the knowledge internalization stage, thereby reducing hallucinations and improving interpretability and robustness.
@@ 2 Related Work
Visual Mathematical Reasoning with MLLMs. Solving visual mathematical problems (e.g., geometry diagrams, algebraic plots, and etc) requires both strong reasoning ability and accurate interpretation of visual primitives and symbolic content (Yan et al., 2025; Qiao et al., 2025b; Zhang et al., 2025b). Most previous works are dedicated to improving the reasoning process, including chain-of-thought strategies (Xu et al., 2025a; Deng et al., 2024), tool-aided reasoning (Trinh et al., 2024; Chen et al., 2023b), test time scaling (Wang et al., 2025e; Hosseini et al., 2024), and reinforcement learning (Wang et al., 2025c; Jiang et al., 2025). Several recent works (Guo et al., 2025c; Jia et al., 2024b; Wei et al., 2024) suggest that one of the major bottlenecks in visual mathematical reasoning is inaccurate visual comprehension. They typically decouple perception from reasoning, and strengthen perception either by designing specialized visual encoders (Zhang et al., 2025a) or by introducing auxiliary visual tasks (Chen et al., 2025a). However, prior works ignore a key issue of whether correctly extracted visual cues are indeed faithfully incorporated into subsequent reasoning.
@@ 3.1 Forming Trustworthy Perception with Synergistic Visual Rewards
We then compute the cost matrix $\mathcal{C}$ between GT Primitives $\mathcal{G}$ and predicted Primitives $\mathcal{P}$ and apply the Hungarian matching algorithm (Kuhn, 1955) to obtain the optimal one-to-one matching $\mathcal{H}$ that minimizes the total cost $\mathcal{S}_{\text{VPR}}$. The VPR offers interpretable, geometry-aware semantic supervision that avoids the pitfalls of pixel-level noise or black-box embedding similarity. More details are provided in Appendix §C.2.
@@ 3.2 Alleviating Reasoning Drift with Knowledge Internalization Reward
Rewarding Faithful Knowledge Internalization. To address these systematic weaknesses, we introduce the Knowledge Internalization Reward (IntlzR), a trained reward model that evaluates whether each reasoning chain remains faithful to the internalized visual representation, thereby enforcing raw perception as the primary substrate of inference (e.g., reasoning drift). Within CogFlow, we train the IntlzR to enhance internalization explicitly, supported by the curated MathCog-IntlzR (see Appendix §C.3 for more details). This dataset is constructed from positive–negative pairs, with each positive matched to five corresponding negatives; positives are sampled from MathCog-SFT, and negatives are synthesized by injecting the five error types described above, providing fine-grained supervision for distinguishing grounded from unanchored reasoning. To better leverage diverse negative signals in IntlzR and adaptively emphasize the most challenging trajectories while improving gradient efficiency and training stability, we adopt Softmax-DPO (Chen et al., 2024a) for optimization:
@@ 3.3 Strengthening Multi-Step Visual Reasoning with VGPO
Then, the training optimization objective $\mathcal{L}$ is formulated as:
@@ 4.1 Experimental Setup
Baselines. We compare our method with a wide range of closed-source MLLMs, such as GPT series (OpenAI, 2024; 2023; 2025), Gemini series (Team et al., 2025), DouBao-1.5-pro (Guo et al., 2025b), GLM-4.5V (Hong et al., 2025), Claude (Anthropic, 2024) and Qwen-VL-Plus (Bai et al., 2023), as well as open-source MLLMs, including InfiMM-Math (Han et al., 2025), InternVL2.5 (Chen et al., 2024b), InternLM-XC2 (Dong et al., 2024), Math-LLaVA (Shi et al., 2024), MultiMath-7B (Peng et al., 2024a), SVE-Math (Zhang et al., 2025b), VLM-R1 (Shen et al., 2025), MathFlow (Chen et al., 2025a), DVLR (Guo et al., 2025c), VCAR (Jia et al., 2024b), VL-Rethinker (Wang et al., 2025a), ThinkLite-VL (Wang et al., 2025d) and SPHINX-MoE (Liu et al., 2024a). Since some concurrent works have yet to release their models, we report the results provided in their original papers, some of which do not include category-wise performance.
Implementation Details. We initialize CogFlow with Qwen2.5-VL-7B (Bai et al., 2025) and train it on our curated MathCog dataset. For cold start, we first train CogFlow on MathCog-SFT subset for 2 epochs with a learning rate of $1\times 10^{-5}$ and a batch size of 64. Subsequently, CogFlow is optimized using VGPO on MathCog-RL subset for 1 epoch with a learning rate of $1\times 10^{-6}$ and a batch size of 16. The IntlzR reward model is based on Qwen2.5-VL-3B and trained on MathCog-IntlzR subset for 3 epochs with a learning rate of $7\times 10^{-6}$ and a batch size 64. All our models are trained on 16 NVIDIA A100 GPUs. Please refer to Appendix §C for more details.
@@ 4.2 Main Results
As shown in Tables 14.2, CogFlow achieves 66.0% accuracy on FlowVerse, 53.9% on MathVerse, 76.8% on Mathvista, 64.1% accuracy on WeMath and 46.2% on DynaMath, surpassing all open-source baselines by large margins. In terms of LogicVista, it delivers competitive performance, trailing only GLM-4.IV-9B. CogFlow’s improvements in visual perception are particularly significant, as evidenced by its superior performance on subsets where visual information dominates. For example, on FlowVerse it reaches 42.7% on Vision Dense, 55.6% on Vision Primary, and 61.1% on Vision Centric. On MathVerse, the gains are consistent across Vision Intensive (44.8%), Vision Dominant (42.1%), and Vision Only (25.7%). These results demonstrate that CogFlow yields more accurate parsing of geometric primitives and relations, leading to stronger perception–reasoning integration. CogFlow also delivers higher-end-to-end problem-solving accuracy across all settings. The largest absolute gains appear in visually demanding subsets, where perception-anchored reasoning is essential. This demonstrates that CogFlow not only perceives diagrams more accurately but also reasons over them more effectively. Figure 19 further shows that CogFlow produces more structurally accurate and semantically consistent predictions compared to strong baselines. Please refer to Appendix §D.12 for more detailed configurations.
@@ 4.3 Ablation Studies
Analysis on Knowledge Internalization Reward. As shown in Figure 6, removing any single Error type consistently degrades performance, indicating that each contributes complementary supervision for internalization. The largest drops arise when excluding omission/misbinding primitives or contradicting geometric constraints, highlighting that correctly binding primitives and respecting core geometric constraints are most critical for keeping reasoning tied to perception. Figure 6 compares different training strategies during training. Vanilla DPO (Rafailov et al., 2023) with all categories offers moderate gains, whereas Softmax-DPO achieves the best results, reaching 66.0% / 56.2% on FlowVerse. This suggests that fine-grained, within-problem preference modeling better captures the nuances of perception-grounded reasoning.
@@
Error Type Analysis. As shown in Figure 9, we employ GPT-5 (OpenAI, 2025) to classify each response on the FlowVerse benchmark into one of four categories: Perception Error, Knowledge Internalization Error, Reasoning Error, and Correct. The first row covers specialized visual-math MLLMs (i.e., MultiMath-7B (Peng et al., 2024a) and SVE-Math-7B (Zhang et al., 2025b)), an GRPO-style model (i.e., VLM-R1 (Shen et al., 2025)), the decoupled method (i.e., MathFlow-7B (Chen et al., 2025a)), and GPT-4o (OpenAI, 2024); the second row traces our variants from the SFT+GRPO baseline through +SynVRs, +IntlzR, and +VGPO to CogFlow. We observe: (1) GPT-4o’s strong overall performance is driven by superior reasoning, yet its diagram perception lags behind specialized 7B models, indicating that strengthening perception and internalization remains essential; (2) although methods such as MathFlow-7B and VLM-R1 reduce Perception Error and Reasoning Error to some extent, these methods do not meaningfully alleviate Knowledge Internalization Error; (3) +SynVRs chiefly reduces Perception Errors ($2\%$), confirming that geometry-aware and semantic visual rewards improve perceptual fidelity; (4) +IntlzR primarily mitigates reasoning drift, reducing Knowledge-Internalization Error by $2\%$ relative to the baseline; and (5) when all components are combined, CogFlow minimizes every error type while maximizing the proportion of correct predictions. Furthermore, we provide a systematic case study in Appendix §D.13 to analyze the error compensation mechanism.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62402432) and in part by the Fundamental Research Funds for the Central Universities (226-202500080).
@@ B.4 Dataset Examples
Specifically, the perception trajectory translates raw diagrammatic pixels into structured geometric primitives (Points, Lines, and Circles) with precise coordinates in a visual parameterized space. For instance, in the leftmost example involving a rectangle, the model identifies specific coordinates for vertices $A$ and $B$ (e.g., A:(7.56,3.5)A:(-7.56,3.5)) and the intersection point $O$. Similarly, in the center and right examples, the model explicitly parameterizes circles by identifying center coordinates and radii, e.g., Circle: (2.52,0.48,7.84)(2.52,-0.48,7.84).
[…]